OpenStudy (anonymous):

find displacement from \[0\le t \le 4\] a(t)=t-5 v(0)=4

OpenStudy (anonymous):

\[s''(t)= a(t)\] so\[s(t)=\int\limits \int\limits a(t)\]

OpenStudy (anonymous):

\[ a(t)=t-5\\ v(t)=\int a(t)dt=\int(t-5)dt\\ v(t)={t^2\over2}-5t+c\\ \text{from the fact} \;v(0)=4,c=4\\ v(t)={t^2\over2}-5t+4\\ s=\int_0^4v(t)dt\\ s=\int_0^4\left({t^2\over2}-5t+4\right)dt \]

OpenStudy (anonymous):

now i do the integral of this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Don't I have to take it between two points as well? \[\int\limits_{0}^{.876}\int\limits_{.876}^{4}\]

OpenStudy (anonymous):

??? where did you get the 0.876 from?

OpenStudy (anonymous):

and the two integrals?

OpenStudy (anonymous):

nvm idk if i'm even asking the question right.